OpenStudy (anonymous):

let f(x)=x^3+1 a) Find f ' (x) b) Simplify f(x+h)-f(x)/h Please dont just tell me out to do it but show me the steps as I do not know how to continue.

OpenStudy (popitree):

3x^2

OpenStudy (anonymous):

@popitree what about part b and ty so much

OpenStudy (popitree):

f'(x) = Lt h->0(f(x+h) - f(x))/h here, {(x+h)^3+1 - (x^3+1)}/h =(x^3+3x^2h+3xh^2+h^3+1 - x^3 - 1)/h =3x^2+(3xh+h^2) now when h tends to 0 the second part vanishes and it will give u derivative

OpenStudy (anonymous):

If i could give you more medals i would thanks man @popitree